OpenStudy (anonymous):

Use inverse operations to write the inverse of the function. f(x) = x + 3 f^-1(x) = ?

hero (hero):

1. Replace f(x) with y 2. Swap x and y 3. Isolate y 4. Replace y with f^-1(x)

hero (hero):

y = x + 3 x = y + 3 x - 3 = y or y = x - 3 f^-1(x) = x - 3

OpenStudy (anonymous):

f(x) = 4x f^-1(x) = ?

OpenStudy (anonymous):

@hero? How would I do that one?